Each of the following temperature ranges is in degrees Fahrenheit. Use the formula F=9/5C+32 to find the corresponding temperature range in degrees Celsius (using interval notation and assume inclusion).
a.below 23∘
b.above 41∘
c.below 14∘
d.above 32∘


Answer by Alan3354(69443)

Convert them to Celsius.
=================
a.below 23º F
C = (F-32)*5/9 = -9*5/9 = -5º C
--> < -5º C
========================
Do the same way.
b.above 41∘
c.below 14∘
d.above 32∘